Assume the total cost of a university education will be $210,000 when your child enters university in 8 years. You presently have $180,000 to invest. What annual rate of interest (APR) must you earn, if the interest is compounded monthly? Enter your answer as a percentage. Do not include the percentage sign in your answer.

Enter your response below rounded to two decimal places.
Answer #1

We use the formula:  
A=P(1+r/12)^12n
where
A=future value
P=present value  
r=rate of interest
n=time period.

210,000=180,000*(1+r/12)^(12*8)

(210,000/180,000)^(1/96)=(1+r/12)

(1+r/12)=1.00160703

r=(1.00160703-1)*12

=1.93(Approx)